the larger of two numbers is 24 less than the square of the smaller. If the larger number is divided by the smaller one, the quotient is -5. Find the larger of the two numbers.
Answer by jorel1380(3719)

Let n be your small number. Then the larger number is nē-24. So:
nē-24/n=-5
nē-24=-5n
nē+5n-24=0
(n+8)(n-3)=0
n=-8 or 3
The larger number is either 40 or -15. ☺☺☺☺